# XLB: A Differentiable Massively Parallel Lattice Boltzmann Library in Python

### v2

@@ 5.2 Multi-GPU Scaling
Strong scaling performance is shown in Figure 10. For this test, the problem size remains fixed at dimensions of $512\times 512\times 512$ while the number of GPUs is increased. Similar to the weak scaling tests, the strong scaling results also indicate remarkable efficiency, with close to 90% efficiency using 8 GPUs.
Figure 9: Weak scaling of XLB library on a DGX A100 multi-GPU setup. Domain sizes are adjusted in proportion to the number of GPUs used (shown with annotations).
Figure 10: Strong scaling of XLB library with fixed problem size on DGX A100 multi-GPU setup.
@@ 5.3 Distributed Scaling
1024
1
8
11448
1296
2
16
21159
1632
4
32
38055
2048
8
64
63987
2688
16
128
105525
3328
32
256
161108
4096
64
512
220332
Table 1: Weak scaling performance of the XLB library on a cluster of DGX A100 nodes, each with 8 GPUs. The test extends up to 64 nodes (512 GPUs), with each JAX process running on a separate node using MPI.
@@ 6 Physics-Based Machine Learning
XLB’s integration with JAX is central to its effectiveness in scientific machine learning, as it enables a unified computational graph for both forward and backward computations. Without JAX, linking disconnected computational graphs of simulations with machine learning models created by a different library would demand substantial engineering, and may potentially result in reduced performance. XLB stands out by allowing for seamless forward and backward propagation across both simulation and neural network components, offering the flexibility to integrate machine learning models at various stages of the simulation workflow, as evidenced in the following demos using the Flax library flax2020github .
@@ 6.1 Enhancing Coarse-grained Simulations Using Deep Learning Correctors
Drawing on the methodology introduced in um2020solver , this example aims to improve the accuracy of a low-fidelity coarse simulation by applying a neural network corrector. We employ the same example of unsteady flow over a cylinder that was discussed in Section 4.3. The corrector is designed to act as a body force that attempts to modify the coarse simulation to resemble the reference solution of the same setup but at a higher resolution (based on an $L^{2}$ norm).
@@
$\{\mathbf{f}_{h}(x_{h},t_{0}),\,\mathbf{f}_{h}(x_{h},t_{0}+\Delta t_{h}),\cdots,\mathbf{f}_{h}(x_{h},t_{0}+k\Delta t_{h})\}$
in which $\mathbf{f}_{h}(x,t)$ corresponds with the distribution function field in all $q$ directions. Here $k$ temporal snapshots are included in the time-series. The above reference solution may be coarse-grained at a lower resolution by using bi-cubic spatial interpolation (available in utils.py) such that $\Delta x_{h}/\Delta x_{l}=\Delta t_{l}/\Delta t_{h}=r>1$. We denote the resulting time-series by,
$\{\mathbf{f}_{l}(x_{l},t_{0}),\mathbf{f}_{l}(x_{l},t_{0}+\Delta t_{l}),\cdots,\mathbf{f}_{l}(x_{l},t_{0}+k\Delta t_{l})\}$
in which $x_{l}$ corresponds with a coarse discretization of the domain. Note that the subscript $h$ and $l$ designate high and low resolutions, respectively. Furthermore, we may define $\mathbf{g}(x_{l},t)$ as the solution of the LBM equations on the coarse grid without external force; a solution that is completely oblivious to the reference solutions $\mathbf{f}_{h}(x_{h},t)$ or its coarse-grained alternative $\mathbf{f}_{l}(x_{l},t)$. The resulting time-series may be written similarly as,
$\{\mathbf{g}(x_{l},t_{0}),\mathbf{g}(x_{l},t_{0}+\Delta t_{l}),\cdots,\mathbf{g}(x_{l},t_{0}+k\Delta t_{l})\}$
@@
$\{\mathbf{g}^{nn}(x_{l},t_{0}),\mathbf{g}^{nn}(x_{l},t_{0}+\Delta t_{l}),\cdots,\mathbf{g}^{nn}(x_{l},t_{0}+k\Delta t_{l})\}$
The latter is obtained after consecutive LBM steps, each including a sequence of operations that may be denoted collectively by $\mathscr{P}$ which represents collision, applying force, streaming, and applying boundary conditions altogether in order to advance the result one step in time. The deep learning corrector force, $\mathcal{F}^{nn}$, is then developed to minimize the difference between the resulting macroscopic velocity fields obtained based on the down-sampled coarse-grained solution $\{\mathbf{f}_{l}(x_{l},t)\}$ and the corrected solution $\{\mathbf{g}^{nn}(x_{l},t)\}$. This is achieved by minimizing the following loss function $\mathcal{L}$, expressed as:
$\mathcal{L}=$ $\sum_{m}^{m+s}\left[\mathscr{U}_{x}\big{(}\mathbf{g}^{nn}(x_{l},t_{0}+m\Delta t_{l})\big{)}-\mathscr{U}_{x}(\mathbf{f}_{l}(x_{l},t_{0}+m\Delta t_{l}))\right]^{2}$
$+\sum_{m}^{m+s}\left[\mathscr{U}_{y}(\mathbf{g}^{nn}(x_{l},t_{0}+m\Delta t_{l}))-\mathscr{U}_{y}(\mathbf{f}_{l}(x_{l},t_{0}+m\Delta t_{l}))\right]^{2}$ (14)
@@
The neural network adjusts the fluid flow dynamics by exerting the corrector force $\mathcal{F}^{nn}(x,t)$. Using the exact-difference method of kupershtokh2004 , the post-collision distribution functions are adjusted as,
$f_{i}^{\ast}(x,t)=\mathscr{C}(f_{i}(x,t))+\frac{1}{\Delta t}\left[f_{i}^{eq}(\rho,u+\varepsilon\mathcal{F}^{nn})-f_{i}^{eq}(\rho,u)\right],$ (15)
@@
For training, we adopted a reference simulation at high resolution with $n_{x}\times n_{y}=(456\times 120)$, and down-sampled its solution by a factor of $r=6$ to arrive at the coarse-grained reference solution. The deep learning model utilizes a ResNet architecture with four residual blocks, each with a $3\times 3$ kernel size, developed using the Flax library flax2020github . Training cycles begin with the reference solution being used to initialize $\mathbf{f}_{l}$, setting $s=100$ and $m$ to values within $(0,1,\ldots,200)$. Gradient checkpointing optimizes memory use during backpropagation. As training cycles are independent, we employ batching over time steps with a batch size of 20. Training starts after $t_{0}=10,000$ time steps to allow for the development of unsteady vortex shedding in the wake, continuing for $s$ additional time steps. We utilize the Adam optimizer with a learning rate of $10^{-3}$ and train the model for Reynolds numbers $Re=950,1000,1100$, for 50 epochs each, and testing it at $Re=1050$.
@@ 6.2 Differentiable Flow Control with Deep Learning
In our experiment, we arbitrarily set $k=200$. The simulation is based on a 2D square with periodic boundaries on all sides and a resolution of $300\times 300$. The deep learning architecture employed is a feedforward multi-layer perceptron consisting of three layers with 32, 64, and 32 neurons, respectively. The neural network takes in the initial density and perturbs it. The Adam optimizer, with a learning rate of $10^{-3}$, is utilized to minimize the loss function across 300 epochs of training.
@@
The results obtained after solving the above optimization problem are illustrated in Figures 13 and 14. Aside from the qualitatively satisfactory outcome of this demo, this experiment demonstrates how automatic-differentiation embedded in the design of XLB library can be easily harnessed to solve more complex inverse problems in CFD.
@@ 7 Conclusion and Future Work
As we look to the future, the XLB library is set to evolve significantly. The roadmap for development is rich with ambitious features, including but not limited to adjoint-based optimization, data assimilation, machine learning-based computational acceleration, different backends to achieve state-of-the-art performance, and out-of-core computing strategies to handle even larger datasets and simulations. These enhancements, among many others, are currently in various stages of planning and development.
@@ Acknowledgments
The authors wish to express their gratitude to the NVIDIA JAX team for their invaluable support and for generously providing the computational resources essential for performance testing of XLB on distributed GPU clusters. We would like to especially acknowledge and appreciate several helpful tips and comments by Frédéric Bastien at NVIDIA who also oversaw the distributed scaling results presented in Section 5. We would also like to acknowledge Massimiliano Meneghin, Ahmed Mahmoud, Olli Lupton, and Santosh Bhavani for their valuable review of the final manuscript.

### v3

@@ 5.2 Multi-GPU scaling
Strong scaling performance is shown in Figure 11. For this test, the problem size remains fixed at dimensions of $512\times 512\times 512$ while the number of GPUs is increased. Similar to the weak scaling tests, the strong scaling results also indicate remarkable efficiency, with close to 90% efficiency using 8 GPUs. We acknowledge that these multi-GPU scaling results might be accentuated due to the relatively low single-GPU performance of our JAX implementation.
Figure 10: Weak scaling of XLB library on a DGX A100 multi-GPU setup (single precision). Domain sizes are adjusted in proportion to the number of GPUs used (shown with annotations).
Figure 11: Strong scaling of XLB library with fixed problem size on DGX A100 multi-GPU setup (single precision).
@@ 5.3 Distributed scaling
1024 1 8 11448
1296 2 16 21159
1632 4 32 38055
2048 8 64 63987
2688 16 128 105525
3328 32 256 161108
4096 64 512 220332
Table 1: Weak scaling performance of the XLB library on a cluster of DGX A100 nodes, each with 8 GPUs (single precision). The test extends up to 64 nodes (512 GPUs), with each JAX process running on a separate node using MPI.
@@ 6 Physics-based machine learning
XLB’s power in scientific machine learning stems from its native differentiability that is provided through integration with JAX, and hence enables a unified computational graph for both forward and backward computations. Linking simulation packages and machine learning models from different libraries often requires complex integration and can lead to performance losses. XLB stands out by allowing for seamless forward and backward propagation across both simulation and neural network components, offering the flexibility to integrate machine learning models at various stages of the simulation workflow, as evidenced in the following demos using the Flax library [72].
@@ 6.1 Enhancing coarse-grained simulations using deep learning correctors
Drawing on the methodology introduced in [42], this example aims to improve the accuracy of a low-fidelity coarse simulation by applying a neural network corrector. We employ the same example of unsteady flow over a cylinder that was discussed in Section 4.3. The corrector is designed to act as a body force that attempts to modify the coarse simulation to resemble the reference solution of the same setup but at a higher resolution (based on an $L^{2}$ norm).
@@
$\{\bm{f}_{h}(\bm{x}_{h},t_{0}),\,\bm{f}_{h}(\bm{x}_{h},t_{0}+\Delta t_{h}),\cdots,\bm{f}_{h}(\bm{x}_{h},t_{0}+k\Delta t_{h})\}$
in which $\bm{f}_{h}(\bm{x},t)$ corresponds with the distribution function field in all $q$ directions. Here $k$ temporal snapshots are included in the time-series. The above reference solution may be coarse-grained at a lower resolution by using bi-cubic spatial interpolation. We denote the resulting time-series by,
$\{\bm{f}_{l}(\bm{x}_{l},t_{0}),\bm{f}_{l}(\bm{x}_{l},t_{0}+\Delta t_{l}),\cdots,\bm{f}_{l}(\bm{x}_{l},t_{0}+k\Delta t_{l})\}$
in which $\bm{x}_{l}$ corresponds with a coarse discretization of the domain. Note that the subscript $h$ and $l$ designate high and low resolutions, respectively. To ensure that the simulations remain identical across different scales, we employed acoustic scaling such that $\Delta\bm{x}_{h}/\Delta\bm{x}_{l}=\Delta t_{h}/\Delta t_{l}=r<1$.
Furthermore, we may define $\bm{g}(\bm{x}_{l},t)$ as the solution of the LBM equations on the coarse grid without external force; a solution that is completely oblivious to the reference solutions $\bm{f}_{h}(\bm{x}_{h},t)$ or its coarse-grained alternative $\bm{f}_{l}(\bm{x}_{l},t)$. The resulting time-series may be written similarly as,
$\{\bm{g}(\bm{x}_{l},t_{0}),\bm{g}(\bm{x}_{l},t_{0}+\Delta t_{l}),\cdots,\bm{g}(\bm{x}_{l},t_{0}+k\Delta t_{l})\}$
@@
$\{\bm{g}^{nn}(\bm{x}_{l},t_{0}),\bm{g}^{nn}(\bm{x}_{l},t_{0}+\Delta t_{l}),\cdots,\bm{g}^{nn}(\bm{x}_{l},t_{0}+k\Delta t_{l})\}$
The latter is obtained after consecutive LBM steps, each including a sequence of operations that may be denoted collectively by $\mathscr{P}$ which represents collision, applying force, streaming, and applying boundary conditions altogether in order to advance the result one step in time. The deep learning corrector force, $\mathcal{F}^{nn}$, is then developed to minimize the difference between the resulting macroscopic velocity fields obtained based on the down-sampled coarse-grained solution $\{\bm{f}_{l}(\bm{x}_{l},t)\}$ and the corrected solution $\{\bm{g}^{nn}(\bm{x}_{l},t)\}$. This is achieved by minimizing the following loss function $\mathcal{L}$, expressed as:
$\mathcal{L}=$ $\sum_{m}^{m+s}\left[\mathscr{U}_{x}\big{(}\bm{g}^{nn}(\bm{x}_{l},t_{0}+m\Delta t_{l})\big{)}-\mathscr{U}_{x}(\bm{f}_{l}(\bm{x}_{l},t_{0}+m\Delta t_{l}))\right]^{2}$
$+\sum_{m}^{m+s}\left[\mathscr{U}_{y}(\bm{g}^{nn}(\bm{x}_{l},t_{0}+m\Delta t_{l}))-\mathscr{U}_{y}(\bm{f}_{l}(\bm{x}_{l},t_{0}+m\Delta t_{l}))\right]^{2}$ (14)
@@
The neural network adjusts the fluid flow dynamics by exerting the corrector force $\mathcal{F}^{nn}(\bm{x},t)$. Using the exact-difference method of [89], the force term is applied through perturbations in the equilibrium distribution function due to a change in the velocity field $\Delta\bm{u}^{nn}$ such that $\Delta\bm{u}^{nn}=\mathcal{F}^{nn}\Delta t/\rho$. More explicitly, in the presence of an external force, this method modifies the post-collision distribution functions, $f_{i}^{\ast}(\bm{x},t)$, as follows:
$f_{i}^{\ast}(\bm{x},t)=\mathscr{C}(f_{i}(\bm{x},t))+\Delta t\left[f_{i}^{eq}(\rho,\bm{u}+\varepsilon\Delta\bm{u}^{nn})-f_{i}^{eq}(\rho,\bm{u})\right],$ (15)
@@
For training, we adopted a reference simulation at high resolution with $n_{x}\times n_{y}=(456\times 120)$, and down-sampled its solution by a factor of $r=6$ to arrive at the coarse-grained reference solution. The deep learning model utilizes a ResNet architecture with four residual blocks, each with a $3\times 3$ kernel size, developed using the Flax library [72]. Training cycles begin with the reference solution being used to initialize $\bm{f}_{l}$, setting $s=100$ and $m$ to values within $(0,1,\ldots,200)$. Gradient checkpointing optimizes memory use during backpropagation. As training cycles are independent, we employ batching over time steps with a batch size of 20. Training starts after $t_{0}=10,000$ time steps to allow for the development of unsteady vortex shedding in the wake, continuing for $s$ additional time steps. We utilize the Adam optimizer with a learning rate of $10^{-3}$ and train the model for Reynolds numbers $Re=950,1000,1100$, for 50 epochs each, and testing it at $Re=1050$.
@@
While promising, these techniques often suffer from generalization issues, and the corrector’s performance may degrade when applied to different flow regimes or geometries. It is important to note that parallel-in-time (PIT) methods [93] offer an alternative approach to enhancing coarse-grained simulations without deep learning. PIT, under certain use cases, may alleviate generalization issues and provide better computational advantages. The work presented here serves as a proof-of-concept for incorporating deep learning models in an LBM simulation, and further research is needed to explore the full potential and limitations of these approaches.
@@ 6.2 Differentiable flow control with deep learning
In our experiment, we arbitrarily set $k=200$. The simulation is based on a 2D square with periodic boundaries on all sides and a resolution of $300\times 300$. We chose a simple feedforward multi-layer perceptron consisting of three layers with 32, 64, and 32 neurons, respectively, as a baseline for this proof-of-concept experiment. The neural network takes in the initial density and perturbs it. The Adam optimizer, with a learning rate of $10^{-3}$, is utilized to minimize the loss function across 300 epochs of training.
@@
The results obtained after solving the above optimization problem are illustrated in Figures 14 and 15. Aside from the qualitatively satisfactory outcome of this demo, this experiment demonstrates as a proof-of-concept how automatic-differentiation embedded in the design of XLB library can be easily used to solve more complex inverse problems in CFD. While the initial density distribution obtained here may be considered as a noise field and hence not physically viable due to small-scale perturbations to $\rho$, the loss function in Eq. 16 could be refined with physically-informed constraints to identify better initial conditions that yield more realistic flow patterns.
@@ 7 Conclusion and future work
As we look to the future, the XLB library is set to evolve significantly. The roadmap for development is rich with ambitious features, including but not limited to adjoint-based optimization, data assimilation, machine learning-based simulation acceleration, different backends and programming models (e.g. based on Warp [91] and Neon [67]) to achieve state-of-the-art performance, and to enable grid refinement (see [94]) and out-of-core computing strategies in order to handle even larger datasets and simulations. These enhancements, among many others, are currently in various stages of planning and development.
@@ Acknowledgments
The authors wish to express their gratitude to the Nvidia JAX team for their invaluable support and for generously providing the computational resources essential for performance testing of XLB on distributed GPU clusters. We would like to especially acknowledge and appreciate several helpful tips and comments by Frédéric Bastien at NVIDIA who also oversaw the distributed scaling results presented in Section 5, and Arun Raman for doing the scaling runs. We would also like to acknowledge Massimiliano Meneghin, Ahmed Mahmoud, Olli Lupton, and Santosh Bhavani for their valuable review of the final manuscript.